OpenStudy (anonymous):

Simplify the following: (2a^3)^-3/(3b^4)^-3

OpenStudy (anonymous):

(3b^4)/(2a^3) To the negative power means 1 over that term to that power. Ex. 2^-4 means 1/(2^4)

OpenStudy (anonymous):

totally lost

OpenStudy (anonymous):

Tell me what you don't understand?

OpenStudy (anonymous):

(2a^3)^-3=1/(2a^3)^3 (3b^4)^-3=1/(3b^4)^3

OpenStudy (anonymous):

Made a mistake the first time sorry. Now take the top term divided by the bottom term. Since they are in fractions, take the first term times the reciprocal of the second term. So the solution would be 1/(2a^3)^3 times (3b^4)^3/1

OpenStudy (anonymous):

I'm sorry, but I really don't understand, maybe it's my fault because I can't work the equation thing to adequately represent the problem